OpenStudy (anonymous):

5(X+1)-2(1-X)=3(3X+1)

OpenStudy (solomonzelman):

just expand everything

OpenStudy (jadeishere):

Basically, just do PEMDAS. Parentheses Exponents (There aren't any in this equation) Multiplication/division (which ever comes first) Addition/subtraction (Again which ever comes first)

OpenStudy (anonymous):

Here it is useful to use distribution in order to make an easier to solve equation, for example: \[5(x+1)=5x+5\]

OpenStudy (anonymous):

Expand it out like Solomon said. 5(X+1)-2(1-X)=3(3X+1) 5(x + 1) = 5x + 5 -2(1-x) = - 2 + 2x 3(3x + 1) = 9x + 3 5x + 5 - 2 + 2x = 9x + 3 Now group all the numbers with x on one side of the = and all the constant numbers on the other side of the. Can you finish or do you need help?